OpenStudy (rachie19):

A 76.00-pound flask of mercury costs $139.00. The density of mercury is 13.534 g/cm3. a) Find the price of one cubic inch of mercury by calculating the following intermediate values.What is the price of one gram of mercury?What is the price of one cubic centimeter of mercury?It takes 3.300 cubic inches of mercury to make one manometer. Find the price of the mercury used to make 13 manometers by first calculating the cost of mercury for one manometer.What is the price of mercury used to make one manometer?What is the price of mercury used to make 13 manometers?

OpenStudy (popeyes):

76 pound has cost of $139 1 pound has cost of $139/76 or, 435.6 gram has cost of $139/76 = $1.83 1 gram has cost of $0.00412 The volume of flask: \[76 \times \frac{ 435.6 }{ 13.534 } = 2446.10cm^3 = 149.212 \ cubic \ inch \] 1 cubic inch has cost of $139 / 149.212 = $0.93156 Now, 1 cubic inch mercury = $0.93156 So, 3.3 cubic inch mercury or 1 manometer = $3.08748 To make 13 manometer: 13 * 3.0748 = $40.13724